General solutions to Tanx +tan4x+tan7x=tanxtan4xtan7x

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
tan a +tan b+tan c=tan a.tan b.tan c iff tan(a+b+c)=0
here a=x, b=4x, c=7x
so tan(x+4x+7x)=0
=>tan(12x)=0
=>12x=n*pi
therefore, x=(n*pi)/12 [Ans]
Hope it helps.
Soumya Ranjan Mohanty.